# Comparing growth of titania and carbonaceous dusty nanoparticles in weakly magnetised capacitively coupled plasmas

### v1

@@ Abstract
This study investigates the growth cycles of titania and carbonaceous dusty nanoparticles in capacitively coupled radiofrequency plasmas, focusing on the impact of weak magnetic field of approximately 480 Gauss. The growth of dusty nanoparticles in these plasmas is cyclic, with particles reaching their maximum size and subsequently moving out of the plasma, followed by the initiation of a new particle growth cycle. We compare the growth of titania and carbonaceous dusty particles in the presence and absence of a magnetic field. The presence of the magnetic field fastens the growth cycle in both types of dusty plasmas. Langmuir probe measurement of the background plasma parameters such as electron temperature and floating potential reveal radial variations in floating potential but not electron temperature. Furthermore, the magnetic field changes the radial variation of floating potential. The magnetic field also influences the spatial distribution of the two dust cloud differently. It is possible that these differences arises due to the material’s different response to the magnetic field. These findings shed light on the intricate dynamics of titania and carbonaceous dusty plasma particle growth, offering insights into the impact of weak magnetization on dust behavior, which have implications for various applications, from space science to semiconductor processing.
  • February 1, 2024
@@ 1 Introduction
A dusty plasma is a four component system consisting of electrons, ions, neutrals, and nanometer (nm) to micrometer ($\mu$m) sized charged solids, i.e., dust. All the charged species interact with each other and the background neutrals to form a complex, coupled system [1]. Because the dust particles are significantly more massive than the other species, these systems can exhibit a broad range of phenomena such as self-organization [2, 3, 4, 5], clustering [6, 7, 8], and new types of wave phenomena [9, 10]. Moreover, the large mass and size of the dust particle enables direct visualization of these phenomena in the laboratory [11, 12, 13].
@@
Dusty plasmas can be formed in a variety of natural [14, 15, 16, 17, 18], and laboratory environments [19, 20, 21, 22]. Many dusty plasma experiments are performed by directly introducing the dust particles into the plasma. However, if a reactive gas is used to generate the plasma, chemical processes can lead to the spontaneous nucleation and growth of dust particles directly from the gas phase. Studies regarding the formation of dusty plasma have a broad range of applications from understanding the formation of dust in space [23] to semiconductor processing [24]. The majority of the visible universe is made up of plasma and dusty plasmas are believed to be prevalent in planetary rings and star formation [1, 25, 15]. For laboratory studies involving the growth of dusty plasma, the background background plasma is usually ignited from an inert gas such as argon (Ar). Over the past three decades, different kinds of particles growth have been reported in literature. Silicate dusts has been grown in Ar/silane $(SiH_{4})$ plasmas [26]. Carbonaceous dusts [27] and carbon nitride [28] have been grown in Ar/acetylene $(C_{2}H_{2})$ and Ar/$C_{2}H_{2}$/nitrogen plasmas respectively. Polyaniline, which are conducting polymers, have been grown in Ar/aniline $(C_{6}H_{7}N)$ plasmas [29]. Organosilicon been grown in Ar/hexamethyldisiloxane (HDMSO) $\left(C_{6}H_{18}OSi_{2}\right)$ plasmas [30]; similar precursors have been used to grow low dielectric constant thin films in plasma enhanced chemical vapor deposition (PECVD) [31, 32]. Aluminum nanocrystals have been grown in Ar/hydrogen/aluminum trichloride $\left(AlCl_{3}\right)$ plasmas [33]. Titanium dioxide $(TiO_{2})$, also known as titania, nanoparticles have been grown in an Ar/titanium tetraisopropoxide (TTIP) $\left(C_{12}H_{28}O_{4}Ti\right)$ plasmas [34]. Whereas studies for the past 30 years have mostly focused on the growth of carbonaceous and silicate particles in dusty plasma, there has been a shift towards exploring other particles growth in the past 10 years, as given by the examples above. Therefore, there is a need to study the properties of the new kinds of dusty plasma.
@@
The focus of this paper is to grow titania nanoparticles in a weakly-magnetized dusty plasma environment with an accompanying detailed investigation of the cyclical growth process in this system. This growth will be compared with carbonaceous dust growth. Titania dust, along with other oxides of titanium, have been proposed as candidates for the first condensates in primitive solar nebula [35, 36, 37]. In semiconductor processing, TTIP has been commonly used in PECVD to grow $TiO_{2}$ thin films. In general, it is common in PECVD to take at least 10 minutes to nucleate nm to $\mu$m-sized particles prior to thin film growth [38, 39, 40, 41, 42, 43]. Dusty plasma particle growth offers the advantage of nucleating nm-sized particles within milliseconds (ms) [44] and growing to $\geq$ 100 nm within tens of seconds (s) [45].
@@
Figure 1: The dust particle growth processes: (a) Nucleation, (b) Coagulation, and (c) Agglomeration
During growth, the particles form a dust cloud, as shown in Fig. 1(a), that is levitated in the plasma due to a balance of several forces such as gravitational force, electric force, thermophoretic force, ion drag force, and neutral drag force [49, 50, 51]. It is common to have a void, i.e., dust-free region, in the dust cloud. At some time during the agglomeration, the particles accumulate enough mass so that gravitational force becomes dominant and the particles are no longer levitated in the plasma. The particles move away from the plasma and the void expands, as shown in Fig. 1(b). That amount of time from the nucleation until the particles move away from the dust cloud is known as a growth cycle. A new cycle of growth immediately begins in the void, as shown in Fig. 1(c), and similarly reported in other literature [34, 52, 53].
In capacitively coupled radiofrequency (RF) plasmas, particles have grown from different reactive gases, but the timing of their growth cycles have varied between experiments, likely due to differences in experimental set ups. The cycle time in plasma is influenced by various factors, including plasma conditions and the concentration of reactive gases [54]. Nonetheless, a consistent pattern emerges in each case: particles undergo a growth cycle, gradually increasing in size and mass before ultimately losing their electrostatic balance and becoming unconfined. For example, in one Ar/$C_{2}H_{2}$ dusty plasma with 30 cm diameter electrodes, 8 cm electrode spacing, and 10 - 20 W of applied RF power, a cycle length of 35 minutes was measured using Rayleigh/Mie scattering of infrared signals [55]. In another Ar/$C_{2}H_{2}$ dusty plasma with 75 mm diameter electrodes, 25 mm electrode spacing, and 5 - 30 W of applied RF power, a cycle length of 55 s was measured using pictures taken with a USB camera [56]. In yet another experiment using Ar/$C_{2}H_{2}$ dusty plasma, with 10 cm diameter electrodes, 4 cm electrode spacing, and 20 W of RF power, a 2-minute cycle was measured by analyzing the scattered laser beam intensity [57]. Ar/$C_{2}H_{2}$ dusty plasma have been well studied in literature. However, as mentioned above, there has been a shift towards exploring different kinds of materials in plasma. Other gases besides Ar/$C_{2}H_{2}$ have also demonstrated a growth cycle. For example, in the same experimental set up at the previously mentioned one, when switching to Ar/methane plasma, it required 60 W of RF power to initiate plasma ignition and a 15-minute cycle time was measured [57]. Using techniques such as mass spectrometry, optical emission spectroscopy (OES), and measurements of electron temperature and electron density, a 150 s cycle was observed in Ar/HDMSO plasma with a top electrode diameter of 10 cm, bottom electrode diameter of 12 cm, electrode spacing ranging from 3.5 to 4 cm, and an RF power of 30 W [58, 59, 60].
This study reports the growth cycles of nm-sized spherical titania particles in Ar/TTIP plasma and compares them to Ar/$C_{2}H_{2}$ dusty plasma in both magnetized and non-magnetized conditions. Specifically, it investigates the impact of a 480 Gauss magnetic field on cycle times, which can offer precise control over dust size. The growth cycles were monitored using techniques such as OES, laser light scattering and temporal dust size distributions, shedding light on the dynamics of particle formation in these plasmas.
2 Method
Figure 3: (a) Side view showing electrode spacing, laser sheet, and gas lines (Not drawn to scale). (b) Top view showing the laser sent in, and perpendicular measurement by CMOS camera and OES (Not drawn to scale). (c) Side view showing the placement of the permanent magnet rings above and below the top and bottom electrodes respectively. The mid plane between the electrodes at z = 12.5 mm is shown. The coordinate origin (0,0,0) is at the center of the surface of bottom electrode (Not drawn to scale). (d) An photograph of an Ar/TTIP dusty plasma, 10s after plasma ignition. Light from the laser sheet is scattered to show a dust cloud levitated between the electrode.
[…]
Base pressure = 3 $\pm$ 0.3 mTorrOpen TTIP valve or $C_{2}H_{2}$ MFCChamber pressure = 35 $\pm$ 3 mtorrOpen argon MFCChamber pressure = 45 $\pm$ 3 mTorrAdjust isolation valvePressure = 300 $\pm$ 1 mTorr?Laser, CMOS Camera, and OES ONPlasma ONCollect dust when plasma OFFClean chamberYesNo
@@ 2 Method
The chamber had 3 quartz viewing ports. A green (532 nm) laser sheet was used to illuminate the dust cloud in two dimension from one viewing port. A complementary metal-oxide semiconductor (CMOS) camera (xiQ MQ042MG-CM) was used to capture the evolution of the dust cloud during particle growth experiment up to atmost 50 frames per second (FPS) from the second viewing port. OES was used to monitor line intensities between 680 nm and 860 nm from the third viewing port. These are shown in Fig. 2(b). A broadband spectrometer was used (Avantes ULS4096CL). The OES resolution was 0.59 nm, the slit size was 25 $\mu$m, grating was 600 lines/mm, the integration time was 100 ms, and 5 data points were averaged.
When experiments were performed using magnetic field, a pair of axially magnetized permanent magnet rings were used. Each hollow cylindrical magnet was made from neodymium-iron-boron (NdFeB), grade 42 with a 51 mm outer diameter, 25 mm inner diameter, and 6 mm thickness. The magnets were placed above and below the plasma electrodes as shown in Fig. 2(c) and produced a nominal magnetic field at the mid-point between the electrodes of $\approx 480$ Gauss. The magnetic field lines distribution in a two dimensional sheet at the center of the electrode is shown in Fig. 4(a). The coordinate origin (0,0,0) is at the center of the surface of bottom electrode. The field lines were calculated using magpylib library in Python [61, 62]. Moreover, Fig. 4(b) shows the radial profile of $B_{z}$ at the middle of the two electrodes, i.e. z = 12.5 mm, and Fig. 4(c) shows the profile of $B_{z}$ as a function of z at x = 0 mm. The red values are measured using F.W. Bell Gauss/Teslameter Model 5080 and blue lines are calculated using magpylib.
@@ 3 Results
In these particle growth experiments, our goal was to characterize the cyclical formation of titania nanoparticles in the presence and absence of a magnetic field and to compare with the formation of carbonaceous nanoparticles. The growth cycles were assessed through three distinct methods: firstly, by measuring the line intensity of Ar I at 760.5 nm via OES; secondly, by capturing images of laser light scattered from the dust cloud using a CMOS camera; and thirdly, by analyzing the dust size distribution for particles grown at different length of time. Each of these processes is described below as they are applied to this study.
@@ 3.1 Optical Emission Spectroscopy
Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e)
Figure 6: OES of the plasma. No new lines were seen during the presence of either TTIP or $C_{2}H_{2}$ in the plasma. (a) Survey scan between 680 and 860 nm revealing several Ar I peaks. The line at 763.5 nm is an isolated peak as shown in the inset. Its intensity is used to track the particles’ growth cycle. Cyclic variation of the intensity of Ar I (763.5 nm) line: (b) Ar/TTIP dusty plasma without magnetic field $(77\pm 4s)$, (c) Ar/TTIP dusty plasma with magnetic field $(32\pm 3s)$, (d) Ar/$C_{2}H_{2}$ dusty plasma without magnetic field $(115\pm s)$, and (e) Ar/$C_{2}H_{2}$ dusty plasma with magnetic field $(39\pm 1s)$.
The first characterization technique is OES. In spite of the low resolution of the spectrometer, 0.59 nm, we obtained good, time-resolved spectra as shown in Fig. 5(a) whereby several Ar peaks are identified. However, no new lines were seen during to the presence of the reactive gases. This might either be due to the relatively low resolution of the broadband spectrometer or due to the higher flow rate of Ar when compared to reactive gas in the plasma. In particular, while it is observed that all of the emission lines vary as a function of time during the growth process, an examination of the (NIST) database reveals that the argon neutral (Ar I) line at 763.5 nm is a promising candidate for OES analysis. As seen in the inset in Fig. 5(a) - and confirmed by NIST database - this line, which was the electric dipole transition of Ar I from $3s^{2}3p^{5}\left({}^{2}P^{0}_{3/2}\right)4p$ to $3s^{2}3p^{5}\left({}^{2}P^{0}_{3/2}\right)4s$ was a relatively isolated line [63, 64, 65]. Moreover, this line was chosen because it was the highest intensity line according to the survey scan without any other line blending. The cyclic variation of the line was chosen to monitor the growth process of the dusty plasmas. The timeframe of the OES was set from the integration time and number of data averaged. Data was collected roughly every 600 ms. The time when the plasma was turned on was set to zero. For the Ar/TTIP dusty plasma, without the presence of the magnetic field, the cycle time was 77 $\pm$ 4 s, and with the presence of the magnetic field, the cycle time was 32 $\pm$ 3 s, as shown in Fig. 5(b) and 5(c) respectively. For the Ar/$C_{2}H_{2}$ dusty plasma, without the presence of the magnetic field, the cycle time was 115 $\pm$ 5 s, and with the presence of the magnetic field, the cycle time was 39 $\pm$ 1 s, as shown in Fig. 5(d) and 5(e) respectively.

### v4

@@ Abstract
This study compares the growth cycles and spatial distribution of dust cloud for titania and carbonaceous dusty nanoparticles in capacitively coupled radiofrequency plasmas, with and without the presence of a weak magnetic field of approximately 500 Gauss. Findings on cycle time, growth rate, and spatial distribution of dust cloud are discussed. The growth of nanoparticles in these plasmas is cyclic, with particles reaching their maximum size and subsequently moving out of the plasma, followed by the generation of a new particle growth cycle. The presence of the magnetic field speeds up the growth cycle in both plasma. The magnetic field also makes the spatial distribution of the two dust cloud different from each other. Langmuir probe measurement of the background plasma parameters such as electron temperature and floating potential reveal radial variations in floating potential but not electron temperature. Furthermore, the magnetic field changes the radial variation of floating potential. These measurements, however, are not sufficient to explain why the two dust clouds appear differently. It is possible that the differences occur due to a gradient in the radial distribution of the magnetic field.
  • July 17, 2024
@@ 1 Introduction
A dusty plasma is a four component system consisting of electrons, ions, neutrals, and nanometer (nm) to micrometre ($\mu$m) sized charged solids, i.e., dust. All the charged species interact with each other and the background neutrals to form a complex, coupled system [1, 2, 3, 4]. Since the dust particles are significantly more massive than the other species, these systems can exhibit a broad range of phenomena such as self-organization [5, 6, 7, 8], clustering [9, 10, 11, 12], and waves phenomena [13, 14]. Moreover, the large size of the dust particle enables direct visualization of these phenomena in the laboratory [15, 16, 17].
@@
Dusty plasmas can be formed in a variety of natural [18, 19, 20, 21, 22], and laboratory environments [23, 24, 25, 26]. Many dusty plasma experiments are performed by directly introducing the dust particles into the plasma [27, 28, 29]. However, if a reactive gas is used to generate the plasma, chemical processes can lead to the spontaneous nucleation and growth of dust particles directly from the gas phase. The background plasma is usually ignited from an inert gas such as argon (Ar). Over the past three decades, different kinds of particles growth have been reported such as silicates [30], carbonaceous [31, 32, 33], polyaniline [34] and organosilicon [35]. Recently, there has been a shift towards exploring metal-containing nanoparticles such as aluminium [36], titanium [37] and titanium dioxide $(TiO_{2})$, also known as titania [38]. Studies regarding the formation of dusty plasma have a broad range of applications from understanding the formation of dust in space [39] to semiconductor processing [40]. The majority of the visible universe is made up of plasma and dusty plasmas are believed to be prevalent in planetary rings and star formation [1, 41, 19].
@@
The focus of this paper is to grow titania nanoparticles, using titanium tetraisopropoxide (TTIP) $\left(C_{12}H_{28}O_{4}Ti\right)$ as the precursor, in a weakly-magnetized ($\sim$ 500 Gauss) dusty plasma environment with an accompanying detailed investigation of the cyclical growth process. This growth will be compared with carbonaceous dust grown using acetylene $(C_{2}H_{2})$. Titania dust, along with other oxides of titanium, have been proposed as candidates for the first condensates in primitive solar nebula [42, 43, 44]. In semiconductor processing, TTIP has been commonly used in PECVD to grow $TiO_{2}$ thin films. In general, it is common in PECVD to take at least 10 minutes to nucleate nm to $\mu$m-sized particles prior to thin film growth [45, 46, 47, 48, 49, 50, 51]. Dusty plasma particle growth offers the advantage of nucleating nm-sized particles within milliseconds (ms) [52] and growing to $\geq$ 100 nm within tens of seconds (s) [53].
Dusty plasma particle growth happens in three steps, (i) nucleation, (ii) coagulation, and (iii) agglomeration (accretion) [40, 33, 54, 55, 56, 57]. Energetic electrons from the background plasma constantly bombard the gas molecules to create excited species, ions and radicals. During nucleation, clusters are formed spontaneously from chemical reactions between neutrals, and ions that are trapped in the plasma [58]. These clusters have stochastic charge fluctuations [52, 59, 60]. During coagulation, clusters and neutrals collide and react to form bigger particles [61]. As the size of the particles increase, the rate of coagulation reaches a saturation [62]. At this stage, the particles have accumulated a negative charge. During agglomeration, also known as accretion, new growth occurs on the surface of the negatively charged particles due to positive ions, radicals and neutrals colliding on its surface [63, 64, 65]. The particles’ size in the three steps are in the order of 1 nm, 10 nm and 100 nm, respectively. Coagulation and agglomeration are dominated by chemisorption and physisorption respectively [66].
Figure 1: (Schematics not drawn to scale) The particle growth processes from the reactive gas: (i) Nucleation, (ii) Coagulation, and (ii) Agglomeration. After the first generation of growth ends, a new one begins. The particles continuously grow in cycles as long as the plasma is turned on.
During growth, the particles form a dust cloud, as shown in Fig. 2(a), that is levitated in the plasma due to a balance of several forces such as gravitational, electric, thermophoretic, ion drag, and neutral drag [67, 68, 69, 70]. The forces’ magnitudes depend on the radius of the particles, with gravitational being cubic, electric being linear, and the others being quadratic. Initially, when the particles nucleate and start to grow, gravitational force is negligible and electric and ion drag forces are the dominant forces. As the particles further grow in size ($>$ 100 nm), gravitational force also becomes significant [70]. Later during the agglomeration, the particles accumulate enough mass and size so that gravitational and ion drag forces cause the particles to move away from the bulk plasma towards the chamber’s wall. Throughout the growth, it is common to have a void, i.e., dust-free region, in the dust cloud. The void expands when the particles move away, as shown in Fig. 2(b). The amount of time between nucleation and the void expanding is known as a growth cycle. A new cycle of growth immediately begins in the void, as shown in Fig. 2(c), and similarly reported in other studies [38, 71, 72]. The cyclic growth of nanoparticles occurs continually when the plasma is on. In capacitively coupled radiofrequency (RF) plasmas, particles have grown from different reactive gases, but the timing of their growth cycles have varied between experiments, likely due to geometrical differences in experimental set ups, RF power, and the concentration and combinations of different reactive gases [73, 74]. The cycle time has been measured via different techniques such as microwave cavity resonance spectroscopy [70], camera images of the dust cloud [75], Rayleigh/Mie scattering of infrared signals [31], scattered laser beam intensity [73] and optical emission spectroscopy (OES) [76, 77, 78]. Cycle times as low as 40 seconds [75] and as high as 35 minutes [31] have been reported. Nonetheless, a consistent pattern emerges: particles undergo a growth cycle, gradually increasing in size and mass before ultimately losing their force balance and becoming unconfined.
2 Method
Figure 3: (Schematics not drawn to scale) (a) Side view showing chamber configurations and laser sheet. (b) Top view showing the laser, and measurement by camera and OES. (c) Side view showing the placement of the permanent magnet rings above and below the top and bottom electrodes respectively. The mid plane between the electrodes at z = 12.5 mm is shown. The coordinate origin (0,0,0) is at the centre of the surface of bottom electrode. (d) A photograph of an Ar/TTIP dusty plasma. Laser light scattering show a dust cloud between the electrode.
@@ 2 Method
Base pressure = 3 $\pm$ 0.3 mTorrOpen TTIP valve or $C_{2}H_{2}$ MFCChamber pressure = 35 $\pm$ 3 mtorrOpen argon MFCChamber pressure = 45 $\pm$ 3 mTorrAdjust isolation valvePressure = 300 $\pm$ 1 mTorr?Laser, Camera, and OES ONPlasma ONCollect dust when plasma OFFClean chamberYesNo
@@
The chamber had 3 quartz viewing ports. A green laser sheet was used to illuminate the dust cloud in two dimensions (2D) from one viewing port. A complementary metal-oxide semiconductor (CMOS) camera (xiQ MQ042MG-CM) was used to record the evolution of the dust cloud during particle growth experiment up to atmost 50 frames per second (FPS) from the second viewing port. OES was used to monitor line intensities between 680 nm and 860 nm from the third viewing port. A broadband spectrometer was used (Avantes ULS4096CL). The OES resolution was 0.59 nm, the slit size was 25 $\mu$m, grating was 600 lines/mm, the integration time was 100 ms, and 5 data points were averaged. The schematics are shown in Fig. 3(b).
When experiments were performed using magnetic field, a pair of axially magnetized permanent magnet rings were used. They were made from neodymium-iron-boron (NdFeB), of grade 42 with a 51 mm outer diameter, 25 mm inner diameter, and 6 mm thickness. The magnets were placed above and below the plasma electrodes as shown in Fig. 3(c). This produced a nominal magnetic field at the mid-point between the electrodes of $\approx 500$ Gauss. The magnetic field lines distribution in a two-dimensional plane at the center of the electrode is shown in Fig. 5(a). The coordinate origin (0,0,0) is at the center of the surface of bottom electrode. The field lines were calculated using magpylib library in Python [79, 80]. Moreover, Fig. 5(b) shows the radial profile of $B_{z}$ at the middle of the two electrodes, i.e. z = 12.5 mm, and Fig. 5(c) shows the profile of $B_{z}$ as a function of z at x = 0 mm. The red values are measured using F.W. Bell Gauss/Teslameter Model 5080 and blue lines are calculated using magpylib. In each individual plasma, the partial pressure of the precursor gas was constant, at $\sim$ 35 mTorr. We assume that within each experiment, the plasma parameters are not changing due to variation in the initial density of precursors. Any changes that will eventually be detected should be due to the magnetic field.
@@ 3 Results
Material characterisation were not performed on the samples grown. Instead, we rely on knowledge that was previously reported from the same chamber using either Ar/$C_{2}H_{2}$ or Ar/TTIP. Material grown in Ar/$C_{2}H_{2}$ plasma are purely carbonaceous with aromatic signatures [75]. Material grown in Ar/TTIP plasma are initially amorphous, but they eventually crystallize into either anatase or rutile after air-annealing at high temperature [38, 81]. Carbon was present on the particles grown from Ar/TTIP, most likely due to the organic nature of TTIP [82]; however, there was no indication of carbon bonding with titanium. Therefore, we assume that these dusty plasma are amorphous titania with superficial carbon contamination.
The growth cycles were assessed through three distinct methods: firstly, by measuring the line intensity of Ar I at 760.5 nm via OES; secondly, by capturing images of laser light scattered from the dust cloud using a CMOS camera; and thirdly, by analyzing the dust size distribution for particles grown at different length of time. Each of these processes is described below as they are applied to this study.
@@ 3.1 Optical Emission Spectroscopy
Figure 6: OES of the plasma. No new lines were seen during the presence of either TTIP or $C_{2}H_{2}$ in the plasma. (a) Survey scan between 680 and 860 nm revealing several Ar I peaks. The line at 763.5 nm is an isolated peak. Cyclic variation of the intensity of Ar I (763.5 nm) line: Ar/TTIP dusty plasma (b) without magnetic field $(77\pm 4s)$ and (c) with magnetic field $(32\pm 3s)$, Ar/$C_{2}H_{2}$ dusty plasma (d) without magnetic field $(115\pm s)$ and (e) with magnetic field $(39\pm 1s)$.
The OES is obtained in our experiments using a broadband spectrometer. An example of a time-resolved spectra as shown in Fig. 6(a) whereby several Ar peaks are identified. The electron temperature could not be calculated from the line ratio of different Ar emission lines because the spectrometer has relatively low resolution (0.59 nm) causing several lines to blend as seen in Fig. 6a. Furthermore, the intensity measurements of the spectrometer have not been calibrated. Nevertheless, the cyclic variation in Ar line intensity, which mimics the particle growth cycle, is reported. In particular, while it is observed that all of the emission lines vary as a function of time during the growth process, an examination of the (NIST) database reveals that the argon neutral (Ar I) line at 763.5 nm is a promising candidate for OES analysis. This line was chosen because it was the highest intensity line according to the survey scan without any other line blending. As seen in the inset in Fig. 6(a) - and confirmed by NIST database - this line, which was the electric dipole transition of Ar I from $3s^{2}3p^{5}\left({}^{2}P^{0}_{3/2}\right)4p$ to $3s^{2}3p^{5}\left({}^{2}P^{0}_{3/2}\right)4s$ was relatively isolated [83, 84, 85]. When the reactive gas was present, no new lines corresponding to emission of new species were seen. For example, it was anticipated to detect lines corresponding to either carbon and/or hydrogen arising from the reactive gas(es). The spectrometer may have not detected these lines either due its relatively low resolution and/or due to the higher flow rate of Ar increasing its intensity when compared to other species in the plasma [86].
[…]
Literature suggests that instabilities, in the form of filamentary modes and azimuthally rotating voids, may arise towards the end of the growth cycle [55] or from dust particles laying on the surface of the electrodes [90]. Here, the instabilities were not present. Were instabilities to arise, they would have been detected as rapid fluctuations in OES intensity.